OpenStudy (hpaul24):

Find the sum of the infinite series 1/3+4/9+16/27+64/81+... if it exists 16.75 30.55 8.99 none

OpenStudy (sshayer):

no this formula is applicable only if |r|<1 when r=1 or>1 it approaches infinity

OpenStudy (sshayer):

\[r^n \rightarrow \infty if ~r \geq1\]

geerky42 (geerky42):

\[ \begin{aligned} \dfrac13+\dfrac49+\dfrac{16}{27}+\dfrac{64}{81}+\cdots&=\dfrac13\left( 1+\dfrac43+\dfrac{16}9+\dfrac{64}{27}+\cdots\right)\\ &=\dfrac13\sum_{n=0}^\infty\left(\dfrac43\right)^n \end{aligned} \]

geerky42 (geerky42):

\[\dfrac13\sum_{i=0}^n\left(\dfrac43\right)^i\ge\dfrac13\sum_{i=0}^n1\] And clearly \(\displaystyle \lim_{n\to\infty}\dfrac13\sum_{i=0}^n1 = \infty\) Therefore our summation diverges.